# Resolving the Identity Crisis in Text-to-Image Generation

### v2

@@ Abstract
State-of-the-art text-to-image models demonstrate impressive realism but suffer from a persistent identity crisis when generating scenes with multiple humans: producing duplicate faces, merging identities, and miscounting individuals. We present DisCo (Reinforcement with DiverSity Constraints), a novel reinforcement learning framework that directly optimizes identity diversity both within images and across groups of generated samples. DisCo fine-tunes flow-matching models using Group-Relative Policy Optimization (GRPO), guided by a compositional reward that: (i) penalizes facial similarity within images, (ii) discourages identity repetition across samples, (iii) enforces accurate person counts, and (iv) preserves visual fidelity via human preference scores. A single-stage curriculum stabilizes training as prompt complexity increases, requiring no additional annotations. On the DiverseHumans Testset, DisCo achieves 98.6% Unique Face Accuracy and near-perfect Global Identity Spread, outperforming both open-source and proprietary models (e.g., Gemini, GPT-Image) while maintaining competitive perceptual quality. Our results establish cross-sample diversity as a critical axis for resolving identity collapse in generative models, and position DisCo as a scalable, annotation-free solution for multi-human image synthesis.
@@
11footnotetext: Corresponding authors. Equal contribution.22footnotetext: Qualcomm AI Research is an initiative of Qualcomm Technologies, Inc.
@@ 1 Introduction
Text-to-image models have recently achieved impressive realism and controllability, powered by diffusion models [9, 22, 21] and flow-based training schemes such as rectified flow and flow matching [16, 14]. However, when tasked with generating scenes with multiple people, current systems frequently replicate nearly identical faces, conflate identities, or miscount individuals, undermining realism and limiting practical utility. This limitation was recently pointed out in [2]. This is a severe constraint in synthetic data generation for various applications such as training group photo personalization models, consistent character generation and storytelling, narrative media, educational content creation, and simulation environments for social interaction research. As illustrated in Fig. 1, these failures persist even when overall image quality is high, revealing a bottleneck in identity differentiation within and across generations. We term this fundamental issue as the identity crisis.
@@
Existing text-to-image methods rely mainly on generating realistic and aesthetically pleasing humans [11, 3]. These models do not address identity diversity—especially as the number of people and scene complexity increase. We noticed that Reinforcement learning (RL) has been applied to the above models to optimize non-differentiable objectives such as prompt adherence, aesthetics, or human preferences [1, 12, 27], and GRPO-style algorithms have improved stability and sample efficiency for flow-matching models [15, 26]. Additionally, RL has shown the ability to correct problematic behaviors that may be ingrained in large models through limited or biased training data. Hence it can effectively break “bad habits” learned during pre-training. However, no prior approach explicitly optimizes human-identity diversity both within a single image or across groups of generations for the same prompt.
@@
We introduce DisCo: Reinforcement with DiverSity Constraints. It is a novel, sample-efficient RL framework that directly targets identity diversity in multi-human generation, with a particular focus on both inter-sample and cross-sample uniqueness. DisCo explicitly optimizes for diversity not only in a specific image, but also across groups of generated sample. It mitigates identity repetition and promotes broader coverage of the human facial identity space. It fine-tunes flow-matching text-to-image models using Group-Relative Policy Optimization (GRPO) [15, 26], guided by a compositional reward that: (i) penalizes facial similarity within images, (ii) discourages identity recurrence across samples, (iii) enforces accurate person counts, and (iv) preserves text–image alignment via an HPS-style score. Reinforcement learning enables flexible optimization over heterogeneous, non-differentiable objectives—overcoming the rigidity of supervised fine-tuning, which typically requires large annotated datasets. To ensure robustness as prompt complexity increases, DisCo employs a single-stage curriculum that anneals the prompt distribution from simple to diverse multi-human scenarios [13].
@@
Empirically, DisCo sets a new standard for multi-human generation by resolving identity collapse both within and across samples. It significantly reduces identity duplication and improves cross-sample diversity and fidelity across varied prompts and model backbones (e.g., SDXL/SD3.5, FLUX variants, proprietary models), without requiring auxiliary annotations. On DiverseHumans and MultiHuman-TestBench, DisCo consistently improves Count Accuracy, Unique-Faces, and Global Identity Spread, while maintaining perceptual quality (Figs. 24; Tables 1-2).
Contributions
  • Identity- and Count-aware RL for multi-human scenes: We cast multi-human generation as RL fine-tuning with diversity- and count-based rewards computed from facial embeddings, both within images and across groups of generations.
  • Group-wise diversity reward: We introduce a novel group-relative term that penalizes cross-sample identity repetition, encouraging broader identity exploration and improving policy advantage estimation under GRPO.
@@
  • Single-stage curriculum: A lightweight sampling curriculum improves stability and generalization as the number of requested individuals scales.
@@
  • State-of-the-art cross-sample identity diversity with strong quality: DisCo delivers large gains in identity uniqueness and count accuracy across models and prompts, without extra spatial or semantic annotations.
@@ 2 Related Work
A NeurIPS 2025 study [2] discuses the limitations the above methods on the multi-human generation task. They also identify the bias in Human generation by these models, also pointed out by [4]. In their future work section, they observed that current SOTA methods merge identities, repeat faces, or miscount people—the precise error modes DisCo targets (Fig. 1).
@@
RL and preference-optimization have been used to optimize non-differentiable objectives such as prompt faithfulness, aesthetics, and human preferences [1, 12, 27]. In the flow-matching setting, GRPO provides value-free, group-relative variance reduction and KL-controlled updates, with curriculum and multi-objective extensions to improve stability and diversity [15, 26]. In contrast to prior work that largely optimizes faithfulness, DisCo explicitly encodes facial-identity diversity constraints both intra-image and inter-image, paired with an identity-aware curriculum, yielding robust gains in multi-human scenes while maintaining quality.
@@ 3 Method
In this Section, we discuss our proposed DisCo finetuning approach in detail. We begin by establishing the mathematical foundations in Section 3.1. Section 3.2 introduces our proposed compositional reward function. To handle the increasing complexity as the number of people generated grows, Section 3.3 presents a single-stage curriculum learning strategy that gradually transitions from simple to complex multi-person scenarios.
@@ 3.1 Preliminaries
Notation. Let $c$ be a text prompt (conditioning), and $t\in[0,1]$ index the sampling trajectory from noise (t=1t{=}1) to data (t=0t{=}0). The latent image distribution at time $t$ is denoted by $p_{t}(x)$, and the time grid by $\{t_{k}\}_{k=0}^{K}$ with $t_{0}{=}1>\cdots>t_{K}{=}0$. We write $w_{t}$ for a standard $d$-dimensional Wiener process and use $\mathcal{N}(0,I)$ for the standard Gaussian.
Flow matching and rectified flows.
We consider continuous-time normalizing flows trained with flow matching (FM) [14]. Given a data sample $x_{0}\sim\mathcal{X}_{0}$ and noise $x_{1}\sim\mathcal{N}(0,I)$, rectified flow (RF) [16] defines the linear probability path
$x_{t}=(1-t)\,x_{0}+t\,x_{1},\quad t\in[0,1],$ (1)
and trains a velocity field $v_{\theta}(x_{t},t)$ to regress the target velocity $v=x_{1}-x_{0}$. FM yields efficient, deterministic ODE sampling with few steps and high sample quality.
Denoising as an MDP.
We cast iterative sampling as an MDP $\langle\mathcal{S},\mathcal{A},\rho_{0},P,R\rangle$ with state $s_{k}=(c,t_{k},x_{t_{k}})$, action $a_{k}=x_{t_{k+1}}$, deterministic transition $s_{k+1}=(c,t_{k+1},x_{t_{k+1}})$, and initial distribution $\rho_{0}(s_{0})=(p(c),\,\delta_{t_{0}=1},\,\mathcal{N}(0,I))$. The policy is $\pi_{\theta}(a_{k}\mid s_{k})=p_{\theta}(x_{t_{k+1}}\mid x_{t_{k}},c)$, and we compute a terminal reward $R(s_{K})=r(x_{t_{K}},c)$ at $t_{K}{=}0$ [e.g., 1, 27].
Figure 3: DisCo training overview. Our method fine-tunes text-to-image models using Flow-GRPO with a compositional reward. Given a prompt, the model generates a group of images evaluated by four components: (1) Intra-Image Diversity penalizes duplicate identities within images, (2) Group-wise Diversity promotes variation across the group, (3) Count Accuracy enforces correct person count, and (4) HPS Quality ensures prompt alignment and visual fidelity. The combined reward guides GRPO updates to improve identity consistency and diversity.
From ODE to Marginal-Preserving SDE.
We begin with the deterministic sampler defined by the probability-flow ODE:
$\frac{dx_{t}}{dt}=v_{\theta}(x_{t},t),\quad t\in[0,1].$
To enable exploration during RL while preserving marginals $\{p_{t}\}$, we follow Flow-GRPO [15] and replace the ODE with an Itô SDE:
$dx_{t}=f_{\theta}(x_{t},t)\,dt+\sigma(t)\,dw_{t},$ (2)
which matches the same $p_{t}$ as the ODE. The relation between drift terms is:
$v_{\theta}(x,t)=f_{\theta}(x,t)-\tfrac{1}{2}\sigma(t)^{2}\nabla_{x}\log p_{t}(x),$
allowing controlled stochasticity via $\sigma(t)$ and score-based compensation. We use Flow-GRPO’s model-based score approximation; see the Supplementary Material (Section D) for details.
Trajectory Policy and GRPO Objective.
Discretizing (2) over $K$ steps defines the trajectory policy $\pi_{\theta}(\tau\mid c)=\prod_{k}p_{\theta}(x_{t_{k+1}}\mid x_{t_{k}},c)$, with log-probability $\log\pi_{\theta}(\tau\mid c)=\sum_{k}\log p_{\theta}(x_{t_{k+1}}\mid x_{t_{k}},c)$. Returns $r(\tau,c)$ are computed on the final image $x_{t_{K}}$, with gradients backpropagated through all steps [15]. For each prompt $c$, we sample a group $G=\{\tau_{i}\}_{i=1}^{M}$ and compute normalized advantages:
$\tilde{A}_{i}=\frac{r(\tau_{i},c)-\mu_{c}}{\sigma_{c}+\epsilon},$ (3)
where $\mu_{c}=\tfrac{1}{M}\sum_{i=1}^{M}r(\tau_{i},c)$ and $\sigma_{c}^{2}=\tfrac{1}{M}\sum_{i=1}^{M}\big(r(\tau_{i},c)-\mu_{c}\big)^{2}$. We optimize:
$\max_{\theta}$ $\mathbb{E}_{c}\!\Bigg[\frac{1}{M}\sum_{i=1}^{M}\tilde{A}_{i}\log\pi_{\theta}(\tau_{i}\mid c)\Bigg]$ (4)
$\;-\;\beta_{KL}\,\mathbb{E}_{c}\!\Big[\mathrm{KL}\big(\pi_{\theta}(\cdot\mid c)\,\|\,\pi_{\theta_{\text{ref}}}(\cdot\mid c)\big)\Big].$
@@
where $\pi_{\theta_{\mathrm{ref}}}$ is the frozen base model and $\beta_{KL}$ controls drift and reward hacking. For efficiency, we train with fewer denoising steps ($K_{\mathrm{train}}\ll K_{\mathrm{test}}$); full schedule is used at test time. See the Supplementary Material (Section D) for hyperparameters.
@@ 3.2 Reward Signal
We denote the set of embeddings for image $i$ by $F_{i}=\{f_{i,1},\dots,f_{i,m_{i}}\}$. Identity similarity between embeddings $u,v\in\mathbb{R}^{d}$ is computed using cosine similarity $s(u,v)=\frac{u^{\top}v}{\|u\|_{2}\|v\|_{2}}$, which simplifies to $u^{\top}v$ for $\ell_{2}$-normalized vectors. All similarity computations use $s(\cdot,\cdot)$ unless otherwise noted.
@@
Using this reward, we aim to discourage identity repetition across the group $G$ generated for the same prompt $c$. As the reward needs to be assigned per-image and not per-group, we compute the counterfactual “remove-one” statistic for every image $i$. Let $F_{G}=\bigcup_{i=1}^{M}F_{i}$ denote all faces across the group and define
@@
Pseudocode is provided in the Supplementary Material (Section A.1). We observe the model performance generally increases when tuned with $r_{\text{img}}^{d}$ and $r_{\text{grp}}^{d}$. However, this model might be susceptible to reward hacking. The nature of hacking, illustrated in the Supplementary Material (Section E.4), includes “grid” artifacts and generating lesser number of humans. Hence, we propose methods to regularize against them.
@@
Quality/alignment term $r_{\mathrm{img}}^{q}$.
To prevent the “grid” artifacts and facial distortions, we use HPSv3 [17] as a reward. We normalize the HPSv3 score to $[0,1]$:
@@ 3.3 Single-stage Curriculum Learning
with annealing weight $\lambda_{t}=\left(\frac{t}{t_{\text{curriculum}}}\right)^{\gamma_{c}}$, where $\gamma_{c}>1$ controls how long the curriculum remains biased toward simple prompts. This strategy ensures a gradual increase in complexity, transitioning from simple to uniform sampling across all prompt types. For further details, refer to Supplementary Material Sections A.2 and D, which describe the curriculum algorithm and hyperparameter settings, respectively.

### v3

@@ Abstract
State-of-the-art text-to-image models suffer from a persistent identity crisis when generating scenes with multiple humans: producing duplicate faces, merging identities, and miscounting individuals. We present DisCo (Reinforcement with Diversity Constraints), a reinforcement learning framework that directly optimizes identity diversity both within images and across groups of generated samples. DisCo fine-tunes flow-matching models using Group-Relative Policy Optimization (GRPO), guided by a compositional reward that: (i) penalizes facial similarity within images, (ii) discourages identity repetition across samples, (iii) enforces accurate person counts, and (iv) preserves visual fidelity and prompt alignment via human preference scores. A single-stage curriculum stabilizes training as prompt complexity increases. Importantly, this method does not require any real data. On the DiverseHumans Testset, DisCo achieves 98.6% Unique Face Accuracy and near-perfect Global Identity Spread, outperforming open-source and proprietary models (e.g., Gemini, GPT-Image) while maintaining perceptual quality. Our results establish cross-sample diversity as a critical axis for resolving identity collapse, positioning DisCo as a scalable, annotation-free solution for multi-human image synthesis. Project page: https://qualcomm-ai-research.github.io/disco/.
@@
11footnotetext: Corresponding authors. Equal contribution.22footnotetext: Qualcomm AI Research, an initiative of Qualcomm Technologies, Inc.
@@ 1 Introduction
Text-to-image models have recently achieved impressive realism and controllability, powered by diffusion models [11, 27, 25] and flow-based training schemes such as rectified flow and flow matching [20, 18]. However, when tasked with generating scenes with multiple people, current methods frequently replicate nearly identical faces or miscount individuals, undermining realism and limiting practical utility. This limitation was recently pointed out in [3]. This is a severe constraint in synthetic data generation for various applications such as training group photo personalization models, consistent character generation and storytelling, narrative media, educational content creation, and simulation environments for social interaction research. As illustrated in Fig. 1, these failures persist even when overall image quality is high, revealing a bottleneck in identity differentiation within and across generations. We term this fundamental issue as the identity crisis.
@@
Existing text-to-image methods rely mainly on generating realistic and aesthetically pleasing humans [15, 5]. These models do not address identity diversity, especially as the number of people and scene complexity increase. Reinforcement learning (RL) has been applied to the above models to optimize non-differentiable objectives such as prompt adherence, aesthetics, or human preferences [2, 16, 32], and GRPO-style algorithms have improved stability and sample efficiency for flow-matching models [19, 31]. Additionally, RL has shown the ability to correct problematic behaviors that may be ingrained in large models through limited or biased training data. However, no prior approach explicitly optimizes human-identity diversity both within a single image or across groups of generations for the same prompt. Beyond the straightforward application of RL, our work makes several novel observations: (1) applying intra-image diversity rewards alone causes global diversity collapse, as duplicate identities shift across samples rather than being eliminated; (2) diversity optimization introduces reward hacking in the form of face undercounting and grid artifacts, requiring explicit regularization; and (3) specialist models require structured curriculum learning to converge stably on complex multi-human prompts.
@@
We introduce DisCo: Reinforcement with DiverSity Constraints. It is a novel, sample-efficient RL framework that directly targets identity diversity in multi-human generation, with a particular focus on both intra-image and cross-sample uniqueness. DisCo explicitly optimizes for diversity not only within a specific image, but also across groups of generated samples, mitigating identity repetition and promoting broader coverage of human facial identities. It fine-tunes flow-matching text-to-image models using Group-Relative Policy Optimization (GRPO) [19, 31], guided by a compositional reward that: (i) penalizes facial similarity within images, (ii) discourages identity recurrence across samples, (iii) enforces accurate person counts, and (iv) preserves text-image alignment and image quality via an HPS score [21], which we find also reinforces the base model’s ability to follow fine-grained per-person compositional prompts such as clothing and hairstyle specifications. Reinforcement learning enables flexible optimization over heterogeneous, non-differentiable objectives. It overcomes the rigidity of supervised fine-tuning, which typically requires large annotated datasets. To ensure robustness as prompt complexity increases, DisCo employs a single-stage curriculum that anneals the prompt distribution from simple to diverse multi-human scenarios [17].
@@
Empirically, DisCo sets a new standard for multi-human generation by resolving identity collapse both within and across samples. It significantly reduces identity duplication and improves cross-sample diversity and fidelity across varied prompts and model backbones (e.g., SDXL/SD3.5, FLUX variants, proprietary models), without requiring auxiliary annotations. Beyond faces, DisCo’s reward framework generalizes to other visual attributes: training on facial diversity alone improves clothing and body-type variation as a byproduct, and performance on general compositional benchmarks is retained or improved. On DiverseHumans and MultiHuman-TestBench, DisCo consistently improves Count Accuracy, Unique-Faces, and Global Identity Spread, while maintaining perceptual quality.
  • Reinforcement for Diversity: To our knowledge, DisCo is one of the first methods to use reinforcement learning to directly optimize identity diversity in text-to-image generation. We use RL to fine-tune multi-human generation models with identity diversity, quality, prompt-alignment and count-based rewards.
  • Group-wise diversity reward: We uncover a novel finding: reinforcing intra-image diversity alone does not eliminate duplicate identities, as these duplicates silently spread across samples, causing global identity spread to collapse. To address this, we introduce a novel group-relative counterfactual reward (Eq. 6), directly discouraging cross-sample identity repetitions under GRPO.
  • Reward hacking characterizations: Diversity optimization alone turns out to be gameable, as the model learns to generate fewer people and produce unnatural grid-like face arrangements that technically satisfy the reward while looking nothing like a real scene or following the prompt description. We quantify these failure modes and successfully address them through explicit count-control, perceptual quality and prompt-alignment rewards.
@@
  • Single-stage curriculum: We propose a lightweight sampling curriculum which improves stability and generalization as the number of requested individuals scales, and proves particularly critical for specialist models compared to generalist models.
@@
  • State-of-the-art cross-sample identity diversity preserving image quality: DisCo delivers large gains in identity uniqueness and count accuracy across models and prompts, surpassing proprietary systems including GPT-Image-1 on Unique Face Accuracy (98.6% vs. 85.1%) and Global Identity Spread (98.3% vs. 89.8%) without extra spatial or semantic annotations.
@@ 2 Related Work
A NeurIPS 2025 study [3] discusses the limitations the above methods on the multi-human generation task. They also identify the bias in Human generation by these models, also pointed out by [6]. In their future work section, they observed that current SOTA methods merge identities, repeat faces, or miscount people. These are the precise issues DisCo targets (Fig. 1).
@@
RL and preference-optimization have been used to optimize non-differentiable objectives such as prompt faithfulness, aesthetics, and human preferences [2, 16, 32]. In the flow-matching setting, GRPO provides value-free, group-relative variance reduction and KL-controlled updates, with curriculum and multi-objective extensions to improve stability and diversity [19, 31]. In contrast to prior work that largely optimizes faithfulness, DisCo explicitly encodes identity diversity constraints both intra-image and inter-image, paired with an identity-aware curriculum, yielding robust gains in multi-human scenes while maintaining quality. A concurrent work [13] also uses RL for diverse video generation, updating the prompt space of video models.
@@ 3 Method
In this Section, we discuss our proposed DisCo finetuning approach in detail. We establish the mathematical foundations in Section 3.1. Section 3.2 introduces our proposed compositional reward function. Section 3.3 presents a single-stage curriculum learning strategy that gradually transitions from simple to complex multi-person scenarios.
@@ 3.1 Preliminaries
Notation. Let $c$ denote a text prompt, $t\in[0,1]$ index the sampling trajectory from noise (t=1t{=}1) to data (t=0t{=}0), $\{t_{k}\}_{k=0}^{K}$ the discrete time grid with $t_{0}{=}1>\cdots>t_{K}{=}0$, $x_{t_{k}}$ the latent at step $k$, and $p_{t}(x)$ the latent distribution at time $t$.
Flow-GRPO. We follow Flow-GRPO [19] in casting the denoising process as a Markov Decision Process (MDP). Given a text prompt $c$, the state at step $k$ is $s_{k}=(c,t_{k},x_{t_{k}})$, the action is the next latent $a_{k}=x_{t_{k+1}}$, and the policy is $\pi_{\theta}(a_{k}\mid s_{k})=p_{\theta}(x_{t_{k+1}}\mid x_{t_{k}},c)$. A terminal reward $R(s_{K})=r(x_{t_{K}},c)$ is computed on the final image at $t_{K}=0$ [2, 32]. To enable exploration while preserving the marginal distributions $\{p_{t}\}$, the deterministic probability-flow ODE is replaced with a marginal-preserving Itô SDE with controlled stochasticity; we refer the reader to Liu et al. [19] for full derivations. For each prompt $c$, we sample a group $G=\{\tau_{i}\}_{i=1}^{M}$ of $M$ trajectories and compute group-normalized advantages:
$\vskip-5.0pt\tilde{A}_{i}=\frac{r(\tau_{i},c)-\mu_{c}}{\sigma_{c}+\epsilon},\vskip-5.0pt$ (1)
where $\mu_{c}$ and $\sigma_{c}$ are the mean and standard deviation of rewards within the group. The GRPO objective is then:
$\vskip-5.0pt\max_{\theta}$ $\mathbb{E}_{c}\!\Bigg[\frac{1}{M}\sum_{i=1}^{M}\tilde{A}_{i}\log\pi_{\theta}(\tau_{i}\mid c)\Bigg]$ (2)
$\;-\;\beta_{KL}\,\mathbb{E}_{c}\!\Big[\mathrm{KL}\big(\pi_{\theta}(\cdot\mid c)\,\|\,\pi_{\theta_{\text{ref}}}(\cdot\mid c)\big)\Big],$
@@
where $\pi_{\theta_{\mathrm{ref}}}$ is the frozen base model and $\beta_{KL}$ controls drift from the reference policy. For efficiency, we train with fewer denoising steps ($K_{\mathrm{train}}\ll K_{\mathrm{test}}$); the full schedule is used at test time.
Figure 3: DisCo training overview. Our method fine-tunes text-to-image models using Flow-GRPO with a compositional reward. Given a prompt, the model generates a group of images evaluated by four components: (1) Intra-Image Diversity penalizes duplicate identities within images, (2) Group-wise Diversity promotes variation across the group, (3) Count Accuracy enforces correct person count, and (4) HPS Quality ensures prompt alignment and quality. The combined reward guides GRPO updates to improve identity diversity.
@@ 3.2 Reward Signal
We denote the set of embeddings for image $i$ by $F_{i}=\{f_{i,1},\dots,f_{i,m_{i}}\}$. Identity similarity between embeddings $u,v\in\mathbb{R}^{d}$ is computed using cosine similarity $s(u,v)=\frac{u^{\top}v}{\|u\|_{2}\|v\|_{2}}$, which simplifies to $u^{\top}v$ for $\ell_{2}$-normalized vectors. Similarity is denoted by $s(\cdot,\cdot)$ unless otherwise noted.
@@
Using this reward, we aim to discourage identity repetition across the group $G$ generated for the same prompt $c$. As the reward needs to be assigned per-image and not per-group, we compute the counterfactual “remove-one” statistic for every image $i$. Let $F_{G}=\bigcup_{i=1}^{M}F_{i}$ denote all faces across the group. Define
@@
Pseudocode is provided in Appendix A.1. We observe the model performance generally increases when tuned with $r_{\text{img}}^{d}$ and $r_{\text{grp}}^{d}$. However, this model might be susceptible to reward hacking. The nature of hacking, illustrated in Appendix  E.5, includes “grid” artifacts, lesser adherence to prompt and generating lesser number of humans. Hence, we propose methods to regularize against them.
@@
Quality/Alignment $r_{\mathrm{img}}^{q}$.
To prevent grid artifacts, lesser prompt adherence & distorted faces, we use HPSv3 [21] as a reward. We normalize HPSv3 score to $[0,1]$:
@@
Beyond preventing visual artifacts, we find that the HPSv3 reward also preserves and improves the base model’s compositional instruction-following. Despite no explicit training for per-person attribute control, DisCo faithfully follows fine-grained per-person specifications such as clothing style and hairstyle alongside identity diversity (Appendix F.4). Paired with results on Geneval2(Appendix F.2), it suggests that optimizing human preference scores during RL fine-tuning reinforces the base model’s general prompt-following capability.
@@ 3.3 Single-stage Curriculum Learning
with annealing weight $\lambda_{t}=\left(\frac{t}{t_{\text{curriculum}}}\right)^{\gamma_{c}}$, where $\gamma_{c}>1$ controls how long the curriculum remains biased toward simple prompts. This strategy ensures a gradual increase in complexity, transitioning from simple to uniform sampling across all prompt types. For further details, refer to Appendix A.2 and D, which describe the curriculum algorithm and hyperparameter settings, respectively.